OpenStudy (anonymous):

integrate (x+1)/(x^2+1)

OpenStudy (anonymous):

split it up into: \[\int\limits_{}^{}x/(x^2 +1)\]and\[\int\limits_{}^{}1/(x^2 +1)\] the first uses substitution and the second is equal to an inverse trig function. Hope that helps

OpenStudy (mimi_x3):

Break it up: \[\int\limits\frac{x}{x^{2}+1} dx+\int\limits\frac{1}{x^{2}+1} dx\] Then integrate it one by one

OpenStudy (anonymous):

could you show how to integrate the second part? I'm still a little confused.

OpenStudy (mimi_x3):

\[\int\limits\frac{1}{x^{2}+1} dx\] Is an inverse trig function. When you integrate it its \[\tan ^{-1}x +C\]